# ../../Bibliographies/main.bib

### v1

@@ 4.1. Ordinal Terms
[…]
The novelty is that we will have a second cardinal-like term, $\Xi$, but $\Xi$ will have a function sort—that is, we will have new terms of the form $\Xi(\alpha)$. Like $\Omega$, we will interpret $\Xi$ polymorphically, including expressions $\Xi^{(J)}(\alpha)$ for $J$ a natural number. As we will see, $\Xi$’s behavior as a function will lead us to a different definition of $K_{\Xi}$. Relatedly, it will be convenient to include some variables from the beginning.
@@
  • for any $\alpha$ in $\mathrm{OT}_{\Xi}$, $\vartheta_{\Omega}\alpha$ is in $\mathbb{H}$,
  • for any $\alpha$ in $\mathrm{OT}_{\Xi}$, $\vartheta_{\Xi}\alpha$ is in $\mathbb{H}$.
@@
$\vartheta_{\Omega}$ and $\vartheta_{\Xi}$ are the collapsing operations corresponding to $\Omega$ and $\Xi$ respectively.
Our formal cardinalities are more complicated since we now have two cardinalities for each $J$. So we define the set of formal cardinalities to be {}(×$\{-\infty\}\cup\left(\mathbb{Z}\times\{\Xi,\Omega\}\right)$, with the ordering $-\infty<(J+1,\Xi)<(J+1,\Omega)<(J,\Xi)$.
To make sense of the definitions below, it is helpful to keep the following perspective in mind. We always apply definitions beginning with a superscript of $\cdot^{0,\Omega}$, which indicates that the largest cardinal is $\Omega^{(0)}$. When we apply some definition to an ordinal $\alpha$, we proceed inductively; when we consider some subterm $\beta$, it might be inside various applications of $\vartheta_{\Xi}$ and $\vartheta_{\Omega}$, and the superscript in our definition keeps track of our passage through these.
The normal state of affairs would be that $\alpha$ contains a subterm $\vartheta_{\Xi}\alpha^{\prime}$ which in turn might contain a subterm $\vartheta_{\Omega}\alpha^{\prime\prime}$ which might contain a subterm $\vartheta_{\Xi}\alpha^{\prime\prime\prime}$, and so on, alternating applications of $\vartheta_{\Xi}$ and $\vartheta_{\Omega}$. When we pass through a $\vartheta_{\Xi}$, we increment $(J,\Omega)$ to $(J,\Xi)$, and then passing through $\vartheta_{\Omega}$ increments to $(J+1,\Omega)$.
There are two complications to this picture. The first is that we might fail to alternate—perhaps we encounter an expression [^10] We could prohibit such things—they do not occur during the application to cut-elimination, so they are not necessary—just like we could have forbidden $\vartheta_{3}\Omega_{2}$ in $\mathrm{OT}_{\Omega_{\omega}}$, but we have tried to keep the rules for building terms very simple. $\vartheta_{\Xi}\vartheta_{\Xi}\alpha^{\prime}$. We interpret this as having skipped an intervening $\vartheta_{\Omega}$, so the inner $\vartheta_{\Xi}$ should increment from $(J,\Xi)$ directly to $(J+1,\Xi)$.
We express this in the following definition, which describes what the effect of $\vartheta_{\kappa}$ should be on the cardinality.
@@
We define $(J,\kappa)+\kappa^{\prime}=\left\{\begin{array}[]{ll}(J,\kappa^{\prime})&\text{if }\kappa=\Omega\text{ and }\kappa^{\prime}=\Xi\\ (J+1,\kappa^{\prime})&\text{otherwise}\end{array}\right.$.
The second complication is that we do not allow binding to “skip over” a $\Xi$. Formally, if $\Xi^{(J)}(\beta)$ appears as a subterm of $\alpha$ and is not bound by some $\vartheta_{\Xi}$ then any cardinal which appears in $\beta$ and is not bound by a corresponding $\vartheta$ in $\beta$ is also not bound by $\alpha$. To help motivate why, consider a forbidden arrangement like ${\color[rgb]{0,0,1}\vartheta_{\Xi}}\Xi({\color[rgb]{0,0,1}\Xi}(0))$. Following the logic of the alternative approach described in Section 3.4, we should be able to write this in the form $(\vartheta_{\Xi}s(v))(t)$ for some terms $t$ and $s$. As long as we restrict $t$ to be an ordinal, we could not do this.
We build this into the syntax by interpreting $\Xi^{(J)}(\alpha)$ so that everything in $\alpha$ is implicitly already inside $J$ applications of $\vartheta_{\Xi}$. For instance, consider how we will later define substitution; let $\alpha=\vartheta_{\Xi}\vartheta_{\Omega}(v\#\Xi^{(1)}(v))$. Then $\alpha[v\mapsto\Omega^{(0)}]$ will be $\vartheta_{\Xi}\vartheta_{\Omega}(\Omega^{(1)}\#\Xi^{(1)}(\Omega^{(0)}))$—inside $\Xi^{(1)}$ we “already know” we’re inside a level of collapsing, so we use the same notation we would have used outside it.
@@
Let $(J,\kappa)\in\mathbb{Z}\times\{\Xi,\Omega\}$ be an uncountable formal cardinality. We define $\mathbb{FC}^{\geq J,\kappa}(\alpha)$ inductively by:
  • $\mathbb{FC}^{\geq J,\kappa}(\#\{\alpha_{i}\})=\bigcup_{i}\mathbb{FC}^{\geq J,\kappa}(\alpha_{i})$,
  • $\mathbb{FC}^{\geq J,\kappa}(\omega^{\alpha})=\mathbb{FC}^{\geq J,\kappa}(\alpha)$,
  • $\mathbb{FC}^{\geq J,\kappa}(\Omega^{(J^{\prime})})=\left\{\begin{array}[]{ll}\emptyset&\text{if }(J^{\prime},\Omega)<(J,\kappa)\\ \{(J-J^{\prime},\Omega)\}&\text{if }(J,\kappa)\leq(J^{\prime},\Omega)\end{array}\right.$
  • $\mathbb{FC}^{\geq J,\kappa}(\Xi^{(J^{\prime})}(\alpha))=\left\{\begin{array}[]{ll}\mathbb{FC}^{\geq J-J^{\prime},\Omega}(\alpha)&\text{if }(J^{\prime},\Xi)<(J,\kappa)\\ (J-J^{\prime},\Xi)\cup\mathbb{FC}^{\geq J-J^{\prime},\Omega}(\alpha)&\text{if }(J,\kappa)\leq(J^{\prime},\Xi)\end{array}\right.$,
  • $\mathbb{FC}^{\geq J,\kappa}(\vartheta_{\Omega}\alpha)=\mathbb{FC}^{\geq(J,\kappa)+\Omega}\alpha$,
  • $\mathbb{FC}^{\geq J,\kappa}(\vartheta_{\Xi}\alpha)=\mathbb{FC}^{\geq(J,\kappa)+\Xi}\alpha$,
  • $\mathbb{FC}^{\geq J,\kappa}(v^{(J^{\prime})})=\left\{\begin{array}[]{ll}\emptyset&\text{if }(J^{\prime},\Xi)<(J,\kappa)\\ (J-J^{\prime}-1,\Omega)&\text{if }(J,\kappa)\leq(J^{\prime},\Xi)\end{array}\right.$.
We define $\overline{\mathbb{FC}}^{\geq J,\kappa}(\alpha)=\sup\mathbb{FC}^{\geq J,\kappa}(\alpha)$ (where $\sup\emptyset=-\infty$).
The definition above takes $v^{(J)}$ to sit between the formal cardinality $(-J,\Xi)$ and the next lower formal cardinality $(-J-1,\Omega)$.
We next need to define how to shift ordinal terms in a systematic way. The main (but not only) application is that if we have some term $\alpha$ which we want to substitute inside $\beta$; somewhere in a subterm of $\beta$ we may be inside some applications of $\vartheta_{\Xi}$ and $\vartheta_{\Omega}$, so we need to adjust how we refer to things in $\alpha$ so they mean the same thing in a different context. We also need to do the reverse—given a subterm deep inside $\beta$, we want to be able to pull it outside of $\beta$ while preserving the meaning.
@@
We define $\alpha^{\geq J,\kappa}_{\pm J^{\prime},\kappa^{\prime}}$ by:
  • #$\#\{\alpha_{i}\}^{\geq J,\kappa}_{\pm J^{\prime},\kappa^{\prime}}=\#\{(\alpha_{i})^{\geq J,\kappa}_{\pm J^{\prime},\kappa^{\prime}}\}$,
  • $(\omega^{\alpha})^{\geq J,\kappa}_{\pm J^{\prime},\kappa^{\prime}}=\omega^{\alpha^{\geq J,\kappa}_{\pm J^{\prime},\kappa^{\prime}}}$,
  • $(\Omega^{(J^{\prime\prime})})^{\geq J,\kappa}_{\pm J^{\prime},\kappa^{\prime}}=\left\{\begin{array}[]{ll}\Omega^{(J^{\prime\prime})}&\text{if }J<J^{\prime\prime}\\ \Omega^{(J^{\prime\prime}+J^{\prime})}&\text{if }J\geq J^{\prime\prime}\end{array}\right.$,
  • $(\Xi^{(J^{\prime\prime})}(\alpha))^{\geq J,\kappa}_{\pm J^{\prime},\kappa^{\prime}}=\left\{\begin{array}[]{ll}\Xi^{(J^{\prime\prime})}(\alpha^{\geq J-J^{\prime\prime},\Xi}_{\pm J^{\prime},\kappa^{\prime}})&\text{if }(J^{\prime\prime},\Omega)<(J,\kappa)\\ \Xi^{(J^{\prime\prime}+J^{\prime})}(\alpha)&\text{if }(J^{\prime\prime},\Omega)\geq(J,\kappa)\text{ and }\kappa^{\prime}=\Omega\\ \Xi^{(J^{\prime\prime}+J^{\prime}+1)}(\alpha)&\text{if }(J^{\prime\prime},\Omega)\geq(J,\kappa)\text{ and }\kappa^{\prime}=\Xi\end{array}\right.$,
  • $(\vartheta_{\Omega}\alpha)^{\geq J,\kappa}_{\pm J^{\prime},\kappa^{\prime}}=\vartheta_{\Omega}(\alpha^{\geq J,\kappa+\Omega}_{\pm J^{\prime},\kappa^{\prime}})$,
  • $(\vartheta_{\Xi}\alpha)^{\geq J,\kappa}_{\pm J^{\prime},\kappa^{\prime}}=\vartheta_{\Xi}(\alpha^{\geq J,\kappa+\Xi}_{\pm J^{\prime},\kappa^{\prime}})$,
  • $(v^{(J^{\prime\prime})})^{\geq J,\kappa}_{\pm J^{\prime},\kappa^{\prime}}=\left\{\begin{array}[]{ll}v^{(J^{\prime\prime})}&\text{if }(J^{\prime\prime},\Omega)<(J,\kappa)\\ v^{(J^{\prime\prime}+J^{\prime})}&\text{if }(J^{\prime\prime},\Omega)\geq(J,\kappa)\text{ and }\kappa^{\prime}=\Omega\\ v^{(J^{\prime\prime}+J^{\prime}+1)}&\text{if }(J^{\prime\prime},\Omega)\geq(J,\kappa)\text{ and }\kappa^{\prime}=\Xi\end{array}\right.$.
We need to define how substitution will behave before we can define the ordering.
@@
We define $\alpha[v\mapsto^{J,\kappa}\beta]$ by induction on $\alpha$:
  • #$\#\{\alpha_{i}\}[v\mapsto^{J,\kappa}\beta]=\#\{\alpha_{i}[v\mapsto^{J,\kappa}\beta]\}$,
  • $\omega^{\alpha}[v\mapsto^{J,\kappa}\beta]=\omega^{\alpha[v\mapsto^{J,\kappa}\beta]}$,
  • $\Omega^{(J^{\prime})}[v\mapsto^{J,\kappa}\beta]=\Omega^{(J^{\prime\prime})}$,
  • $\Xi^{(J^{\prime})}(\alpha)[v\mapsto^{J,\kappa}\beta]=\Xi^{(J^{\prime})}(\alpha[v\mapsto^{J-J^{\prime},\Omega}\beta])$,
  • $(\vartheta_{\Omega}\alpha)[v\mapsto^{J,\kappa}\beta]=\vartheta_{\Omega}(\alpha[v\mapsto^{J,\kappa+\Omega}]\beta)$,
  • $(\vartheta_{\Xi}\alpha)[v\mapsto^{J,\kappa}\beta]=\vartheta_{\Xi}(\alpha[v\mapsto^{J,\kappa+\Xi}]\beta)$,
  • $w^{(J^{\prime})}[v\mapsto^{J,\kappa}\beta]=\left\{\begin{array}[]{ll}\beta^{\geq 0,\Omega}_{+J,\kappa}&\text{if }v=w\\ w&\text{otherwise}\end{array}\right.$.
As before, we say $v$ is $J,\kappa$-substitutable in $\alpha$ if, in the definition of $\alpha[v\mapsto^{\geq 0,\Omega}\beta]$, $v^{(J^{\prime},\kappa^{\prime})}$ only appears when $J=J^{\prime}$ and $\kappa=\kappa^{\prime}$.
@@
When $\alpha$ is an ordinal term, the unapplication of $\alpha$ is the unique (up to renaming of variables) ordinal term $\bar{\alpha}$ with variables such that:
  • each variable appearing in $\bar{\alpha}$ but not $\alpha$ appears only once,
  • each variable apearing in $\bar{\alpha}$ but not $\alpha$ is $0,\Omega$-substitutable in $\bar{\alpha}$,
  • $\overline{\mathbb{FC}}^{\geq 0,\Omega}(\bar{\alpha})<(0,\Xi)$,
  • $\alpha=\bar{\alpha}[v_{1}\mapsto\beta_{1},\ldots,v_{n}\mapsto\beta_{n}]$ for some sequence $\beta_{1},\ldots,\beta_{n}$ where each $\beta_{i}$ has the form $\Xi^{(0)}(\beta^{\prime}_{i})$.
That is, $\bar{\alpha}$ is the result of going through $\alpha$ inductively and, each time we find a term $\Xi^{(J)}(\beta)$ whose “true” level is 00, replacing that subterm with a fresh free variable.
@@
The definition of $K_{\Omega}\alpha$ is largely unchanged:
  • $K^{\geq J,\kappa}_{\Omega}\#\{\alpha_{i}\}=\bigcup_{i}K^{\geq J,\kappa}_{\Omega}\alpha_{i}$,
  • $K^{\geq J,\kappa}_{\Omega}\omega^{\alpha}=K^{\geq J,\kappa}_{\Omega}\alpha$,
  • $K^{\geq J,\kappa}_{\Omega}\Omega^{(J^{\prime})}=\left\{\begin{array}[]{ll}\{\Omega^{J^{\prime}-J}\}&\text{if }J^{\prime}\geq J\\ \emptyset&\text{if }J^{\prime}<J\end{array}\right.$,
  • $K^{\geq J,\kappa}_{\Omega}\Xi^{(J^{\prime})}(\alpha)=\left\{\begin{array}[]{ll}\{\Xi^{J^{\prime}-J}(\alpha)\}&\text{if }J^{\prime}\geq J\\ K^{\geq J-J^{\prime},\Omega}_{\Omega}\alpha&\text{if }J^{\prime}<J\end{array}\right.$,
@@
  • $K^{\geq J,\kappa}_{\Omega}\vartheta_{\Omega}\alpha=\left\{\begin{array}[]{ll}(\vartheta_{\Omega}\alpha)^{\geq 0,\Omega}_{-J,\kappa}&\text{if }\overline{\mathbb{FC}}^{\geq J,\kappa}(\vartheta_{\Omega}\alpha)<(0,\Omega)\\ K^{\geq J,\kappa+\Omega}_{\Omega}\alpha&\text{if }\overline{\mathbb{FC}}^{\geq J,\kappa}(\vartheta_{\Omega}\alpha)\geq(0,\Omega)\end{array}\right.$,
  • $K^{\geq J,\kappa}_{\Omega}\vartheta_{\Xi}\alpha=\left\{\begin{array}[]{ll}(\vartheta_{\Xi}\alpha)^{\geq 0,\Omega}_{-J,\kappa}&\text{if }\overline{\mathbb{FC}}^{\geq J,\kappa}(\vartheta_{\Xi}\alpha)<(0,\Omega)\\ K^{\geq J,\kappa+\Xi}_{\Omega}\alpha&\text{if }\overline{\mathbb{FC}}^{\geq J,\kappa}(\vartheta_{\Xi}\alpha)\geq(0,\Omega)\end{array}\right.$,
  • $K^{\geq J,\kappa}_{\Omega}v^{(J^{\prime})}=\left\{\begin{array}[]{ll}\{v^{(J^{\prime}-J)}\}&\text{if }J^{\prime}\geq J\\ \emptyset&\text{if }J^{\prime}<J\end{array}\right.$.
The definition of $K^{\geq J,\kappa}_{\Xi}\alpha$ contains a crucial new step:
  • $K^{\geq J,\kappa}_{\Xi}\#\{\alpha_{i}\}=\bigcup_{i}K^{\geq J,\kappa}_{\Xi}\alpha_{i}$,
  • $K^{\geq J,\kappa}_{\Xi}\omega^{\alpha}=K^{\geq J,\kappa}_{\Xi}\alpha$,
  • $K^{\geq J,\kappa}_{\Xi}\Omega^{(J^{\prime})}=\left\{\begin{array}[]{ll}\{\Omega^{J^{\prime}-J}\}&\text{if }J^{\prime}\geq J\\ \emptyset&\text{if }J^{\prime}<J\end{array}\right.$,
  • $K^{\geq J,\kappa}_{\Xi}\Xi^{(J^{\prime})}(\alpha)=\left\{\begin{array}[]{ll}\{\Xi^{J^{\prime}-J}(\alpha)\}&\text{if }J^{\prime}\geq J\\ K^{\geq J-J^{\prime},\Omega}_{\Xi}\alpha&\text{if }J^{\prime}<J\end{array}\right.$,
  • $K^{\geq J,\kappa}_{\Xi}\vartheta_{\Omega}\alpha=\left\{\begin{array}[]{ll}(\vartheta_{\Omega}\alpha)^{\geq 0,\Omega}_{-J,\kappa}&\text{if }\overline{\mathbb{FC}}^{\geq J,\kappa}(\vartheta_{\Omega}\alpha)<(0,\Xi)\\ K_{\Xi}^{\geq J,\kappa+\Xi}\alpha&\text{if }\overline{\mathbb{FC}}^{\geq J,\kappa}(\vartheta_{\Omega}\alpha)\geq(0,\Xi)\end{array}\right.$,
  • $K^{\geq J,\kappa}_{\Xi}\vartheta_{\Xi}\alpha=\left\{\begin{array}[]{ll}\left(\overline{(\vartheta_{\Xi}\alpha)^{\geq 0,\Omega}_{-J,\Xi}}\right)^{\geq 0,\Omega}_{-0,\kappa}&\text{if }\overline{\mathbb{FC}}^{\geq J,\kappa}(\vartheta_{\Xi}\alpha)\leq(0,\Xi)\\ K_{\Xi}^{\geq J,\kappa+\Xi}\alpha&\text{if }\overline{\mathbb{FC}}^{\geq J,\kappa}(\vartheta_{\Xi}\alpha)>(0,\Xi)\end{array}\right.$,
  • $K_{\Xi}^{\geq J,\kappa}v^{(J^{\prime})}=\left\{\begin{array}[]{cc}v^{(J^{\prime}-J)}&\text{if }J^{\prime}\geq J\\ \emptyset&\text{if }J^{\prime}<J\end{array}\right.$.
Note the critical difference in the definition of $K_{\Xi}\vartheta_{\Xi}\alpha$. When we only have $\overline{\mathbb{FC}}^{\geq J,\kappa}(\vartheta_{\Xi}\alpha)\leq(0,\Xi)$, instead of strictly less than. We cannot fully shift by $-J,\kappa$, because this could lead to collissions. So first shift as much as we can, then apply the unapplication operation to replace all instances of $\Xi^{(0)}$ with variables, then finish the shifting.
For instance, consider the ordinal $\vartheta_{\Xi}\alpha={\color[rgb]{1,0,0}\vartheta_{\Xi}}\vartheta_{\Omega}{\color[rgb]{0,0,1}\vartheta_{\Xi}}{\color[rgb]{1,0,0}\Xi}^{(1)}(0)$. When we compare this to other ordinals, we will look at the set $K^{\geq 1,\Omega}_{\Xi}\alpha=\{\vartheta_{\Xi}v\}$.
@@
That is, this ordinal recognizes that the ordinal $\alpha$ includes applying the function $\vartheta_{\Xi}v$ to $\Xi^{(0)}$; in our definition of $<$ below, we will reflect this by requiring that $\vartheta_{\Xi}\alpha$ be closed under the function $\vartheta_{\Xi}v$, in the sense that, for any $\beta<\vartheta_{\Xi}\alpha$, $(\vartheta_{\Xi}v)[v\mapsto\beta]<\vartheta_{\Xi}\alpha$. (Note that $\vartheta_{\Xi}\alpha$ itself is not of this form, because the definition of substitution prevents anything in $\beta$ from getting bound during the substitution.)
@@
This perhaps looks more conventional in the alternative notation described at the end of the previous section—we could think of this ordinal as $\vartheta_{\Xi}\vartheta_{\Omega}(\vartheta_{\Xi}v)(\Xi(0))$, which makes identifying $\vartheta_{\Xi}v$ as an element of $K^{\geq 1}_{\Xi}\alpha$ a natural interpretation of the $K$ operator in this context.
The definition of the order is familiar, but with an addition in the $\vartheta_{\Xi}$ case, and a small complication: we need to consider whether we are comparing in the position where we are immediately above an $\Omega$ cardinal, or above a $\Xi$ cardinal. (To see why this is necessary, observe that we have $\Xi^{(1)}(0)<\Omega^{(1)}<\Xi^{(0)}(0)$, which we will understand to be the $<_{\Omega}$ comparison; if we substitute the second and third terms into $\vartheta_{\Xi}v$, we should also have $\vartheta_{\Xi}\Omega^{(1)}<\vartheta_{\Xi}\Xi^{(1)}(0)$, which leads us to expect that, inside $\vartheta_{\Xi}$, $\Omega^{(1)}<\Xi^{(1)}(0)$; we resolve this by interpreting the setting inside $\vartheta_{\Xi}$ as being the $<_{\Xi}$ comparison.) The only difference between these perspectives is the order between $\Omega^{(J)}$ and $\Xi^{(J)}$ expressions.
@@
We define the ordering $\alpha<_{\kappa}\beta$ by:
  • #$\#\{\alpha_{i}\}<_{\kappa}\#\{\beta_{j}\}$ if there is some $\beta_{j_{0}}\in\{\beta_{j}\}\setminus\{\alpha_{i}\}$ such that, for all $\alpha_{i}\in\{\alpha_{i}\}\setminus\{\beta_{j}\}$, $\alpha_{i}<_{\kappa}\beta_{j_{0}}$,
@@
    • $\beta<_{\kappa}\#\{\alpha_{i}\}$ if there is some $i$ with $\beta\leq_{\kappa}\alpha_{i}$,
    • #$\#\{\alpha_{i}\}<_{\kappa}\beta$ if, for all $i$, $\alpha_{i}<_{\kappa}\beta$,
  • $\omega^{\alpha}<_{\kappa}\omega^{\beta}$ if $\alpha<\beta$,

### v2

@@ 4.1. Ordinal Terms
We next describe a system introducing our new device, replacing $\Omega$ with a cardinal-like term $\Xi$, but $\Xi$ will have a function sort—that is, we will have new terms of the form $\Xi(\alpha)$. Like $\Omega$, we will interpret $\Xi$ polymorphically, including expressions $\Xi^{(J)}(\alpha)$ for $J$ a natural number. Some device like polymorphism is necessary here—we will need to consider terms like $(v(\Xi(0)))[v\mapsto\vartheta w]$, which needs a way to “protect” the $\Xi(0)$ in the argument of $v$ from being collapsed when it is substituted into $\vartheta w$. As we will see, $\Xi$’s behavior as a function will lead us to a different definition of $K_{\Xi}$. Relatedly, it will be convenient to include some variables from the beginning.
@@
  • for any $\alpha$ in $\mathrm{OT}_{\Xi}$, $\vartheta\alpha$ is in $\mathbb{H}$.
@@
Our formal cardinalities are the same as in the previous section.
We can now define operations like shifting and substitution similar to the way we did above. There is one subtlety, which is that in $\Xi^{(J)}(\alpha)$, we wish everything in $\alpha$ to have cardinality at most [^9] Explicitly, we need to prohibit something like ${\color[rgb]{0,0,1}\vartheta}\Xi({\color[rgb]{0,0,1}\Xi}(0))$, where the inner $\Xi$ is bound while the outer one is free. To help motivate why this expression is incorrect, consider the alternate syntax where we replace polymorphism with $\lambda$ expressions. We should be able to write this in the form $(\lambda v.s(v))(t)$ for some terms $s$ and $t$; as long as $t$ is an ordinal, we cannot do this. that of $\Xi$. To enforce this syntactically, we take $\Xi^{(J)}(\Xi^{(0)}(0))$ to mean that the inner $\Xi^{(0)}$ is at the same level as the outer one.
@@
We define $\alpha^{\leq J}_{\pm J^{\prime}}$ by:
  • #$\#\{\alpha_{i}\}^{\leq J}_{\pm J^{\prime}}=\#\{(\alpha_{i})^{\leq J}_{\pm J^{\prime}}\}$,
  • $(\omega^{\alpha})^{\leq J}_{\pm J^{\prime}}=\omega^{\alpha^{\leq J}_{\pm J^{\prime}}}$,
  • $(\Xi^{(J^{\prime\prime})}(\alpha))^{\leq J}_{\pm J^{\prime}}=\left\{\begin{array}[]{ll}\Xi^{(J^{\prime\prime}+J^{\prime})}(\alpha)&\text{if }J^{\prime\prime}\leq J\\ \Xi^{(J^{\prime\prime})}(\alpha^{\leq J-J^{\prime\prime}}_{\pm J^{\prime}})&\text{if }J^{\prime\prime}>J\\ \end{array}\right.$,
  • $(\vartheta\alpha)^{\leq J}_{\pm J^{\prime}}=\vartheta(\alpha^{\leq J}_{\pm J^{\prime}})$,
  • $(v^{(J^{\prime\prime})})^{\leq J}_{\pm J^{\prime},\kappa^{\prime}}=\left\{\begin{array}[]{ll}v^{(J^{\prime\prime}+J^{\prime})}&\text{if }J^{\prime\prime}\leq J\\ v^{(J^{\prime\prime})}&\text{if }J^{\prime\prime}>J\\ \end{array}\right.$.
Note that, for this purpose, $v^{(J)}$ sits in the same cardinality as $\Xi^{(J)}$. (It will lie at the very bottom of this cardinality, so it will in other places behave like cardinality $J-1$ instead.)
Next we can define how substituting for the variables $v^{(J)}$ works.
@@
We define $\alpha[v\mapsto^{J}\beta]$ by induction on $\alpha$:
  • $\Xi^{(J^{\prime})}(\alpha)[v\mapsto^{J}\beta]=\Xi^{(J^{\prime})}(\alpha[v\mapsto^{J-J^{\prime}}\beta])$,
  • $(\vartheta\alpha)[v\mapsto^{J}\beta]=\vartheta(\alpha[v\mapsto^{J-1}]\beta)$,
  • $w^{(J^{\prime})}[v\mapsto^{J}\beta]=\left\{\begin{array}[]{ll}\beta^{\leq 0}_{+J}&\text{if }v=w\\ w&\text{otherwise}\end{array}\right.$.
As before, we only want to substitute into $v$ when it is at the right level at the moment of substitution.
@@
We define when $v$ is $J$-substitutable in $\alpha$ by:
  • $v$ is $J$-substitutable in #$\#\{\alpha_{i}\}$ if $v$ is $J$-substitutable in each $\alpha_{i}$,
  • $v$ is $J$-substitutable in $\omega^{\alpha}$ if $v$ is $J$-substitutable in $\alpha$,
  • $v$ is $J$-substitutable in $\Xi^{(J^{\prime})}(\alpha)$ if either $v$ does not appear in $\alpha$, or $J^{\prime}>J^{\prime}$ and $v$ is $J-J^{\prime}$ substitutable in $\alpha$,
  • $v$ is $J$-substitutable in $(\vartheta\alpha)$ if $v$ is $J-1$-substitutable in $\alpha$,
  • $v$ is $J$-substitutable in $w^{(J^{\prime})}$ if $v\neq w$ or $J=J^{\prime}$.
We next collect up the cardinalities present in an ordinal term.
@@
We define $\mathbb{FC}^{\leq J}(\alpha)$ inductively by:
  • $\mathbb{FC}^{\leq J}(\#\{\alpha_{i}\})=\bigcup_{i}\mathbb{FC}^{\leq J}(\alpha_{i})$,
  • $\mathbb{FC}^{\leq J}(\omega^{\alpha})=\mathbb{FC}^{\leq J}(\alpha)$,
  • $\mathbb{FC}^{\leq J}(\Xi^{(J^{\prime})}(\alpha))=\left\{\begin{array}[]{ll}\{J^{\prime}-J\}\cup\{J^{\prime\prime}+(J^{\prime}-J)\mid J^{\prime\prime}\in\mathbb{FC}^{\leq 0}(\alpha)\}&\text{if }J^{\prime}\leq J\\ \mathbb{FC}^{\leq J-J^{\prime}}(\alpha)&\text{if }J^{\prime}>J\\ \end{array}\right.$,
  • $\mathbb{FC}^{\leq J}(\vartheta\alpha)=\mathbb{FC}^{\leq J-1}\alpha$,
  • $\mathbb{FC}^{\leq J}(v^{(J^{\prime})})=\left\{\begin{array}[]{ll}J^{\prime}-1&\text{if }J^{\prime}\leq J\\ \emptyset&\text{if }J^{\prime}>J\\ \end{array}\right.$.
We define $\overline{\mathbb{FC}}^{\leq J}(\alpha)=\sup\mathbb{FC}^{\leq J}(\alpha)$ (where $\sup\emptyset=-\infty$).
The definition above takes $v^{(J)}$ to sit between the formal cardinality $(J,\Xi)$ and the next lower formal cardinality $(J+1,\Omega)$.
@@
When $\alpha$ is an ordinal term, a parameter of $\alpha$ is an ordinal term $\beta$ in $\mathbb{SC}$ so that there is some $\alpha^{\prime}(v)$ with a fresh 00-substitutable variable $v$ so that with $\alpha=\alpha^{\prime}[v\mapsto^{0}\beta]$.
The canonical abstraction of $\alpha$ is the expresion $\alpha=\bar{\alpha}[v_{1}\mapsto^{0}\beta_{1},\ldots,v_{n}\mapsto^{0}\beta_{n}]$ such that $\overline{\mathbb{FC}}^{\geq 0}(\bar{\alpha})<0$ and each $\beta_{i}$ has the form $\Xi^{(0)}(\beta^{\prime}_{0})$.
That is, the canonical abstraction is the expression we get by pulling out precisely the parameters which are of the form $\Xi^{(0)}(\beta^{\prime}_{0})$.
@@
The definition of $K^{<J}\alpha$ contains a crucial difference:
  • $K^{<J}\#\{\alpha_{i}\}=\bigcup_{i}K^{<J}\alpha_{i}$,
  • $K^{<J}\omega^{\alpha}=K^{<J}\alpha$,
  • $K^{<J}\Xi^{(J^{\prime})}(\alpha)=\left\{\begin{array}[]{ll}\{\Xi^{J^{\prime}-(J-1)}(\alpha)\}&\text{if }J^{\prime}<J\\ K^{\leq J-J^{\prime}}\alpha&\text{if }J^{\prime}\geq J\end{array}\right.$,
@@
  • $K^{<J}\vartheta\alpha=\left\{\begin{array}[]{ll}\left(\overline{(\vartheta\alpha)^{\leq 0}_{-J^{\prime}}}\right)^{\leq 0}_{+1}&\text{if }\overline{\mathbb{FC}}^{\leq 0}(\vartheta\alpha)\leq J\\ K^{\leq J,\kappa+\Xi}\alpha&\text{if }\overline{\mathbb{FC}}^{\leq 0}(\vartheta\alpha)>J\end{array}\right.$,
  • $K^{<J}v^{(J^{\prime})}=\left\{\begin{array}[]{ll}\{v^{(J-(J-1))}\}&\text{if }J^{\prime}<J\\ \emptyset&\text{if }J^{\prime}\geq J\end{array}\right.$.
In the definition of $K\vartheta\alpha$, where we include the case where $\overline{FC}^{\leq 0}(\vartheta\alpha)=J$. In this case, we think of $\vartheta\alpha$ as the result of applying a function to some $\Xi^{(0)}(\beta_{i})$ term, and we place this function in $K^{<J}$.
@@
For instance, consider the ordinal $\vartheta\alpha={\color[rgb]{1,0,0}\vartheta}{\color[rgb]{0,0,1}\vartheta}{\color[rgb]{1,0,0}\Xi}^{(1)}(0)$. When we compare this to other ordinals, we will look at the set $K^{<0}\alpha=\{\vartheta v\}$. That is, this ordinal recognizes that the ordinal $\alpha$ includes applying the function $\vartheta v$ to $\Xi^{(0)}$; in our definition of $<$ below, we will reflect this by arranging that $\vartheta\alpha$ be closed under the function $\vartheta v$, in the sense that, for any $\beta<\vartheta\alpha$, $(\vartheta v)[v\mapsto\beta]<\vartheta\alpha$. (Note that $\vartheta\alpha$ itself is not of this form, because the definition of substitution prevents anything in $\beta$ from getting bound during the substitution.)
@@
This perhaps looks more conventional in the alternative notation—we could think of this ordinal as $\vartheta\left[(\lambda v.\vartheta v)\Xi(0)\right]$, which makes identifying $\lambda v.\vartheta v$ as an element of $K^{<0}\alpha$ a natural interpretation of the $K$ operator in this context.
@@
We define the ordering $\alpha<\beta$:
@@
    • $\beta<\#\{\alpha_{i}\}$ if there is some $i$ with $\beta\leq\alpha_{i}$,
    • #$\#\{\alpha_{i}\}<\beta$ if, for all $i$, $\alpha_{i}<\beta$,
  • $\omega^{\alpha}<\omega^{\beta}$ if $\alpha<\beta$,